# Monocular Lane Detection Based on Deep Learning: A Survey

### v2

@@ VI-D Lane Topology Reasoning
Topology reasoning aims to comprehensively understand road scenes and present drivable routes in autonomous driving. It requires detecting road centerlines and traffic elements, further reasoning their topology relationship, i.e., lane-lane topology, and lane-traffic topology. Directly using vehicle-mounted sensors to detect lane topology has become popular due to their valuable application.
The early topology reasoning is mainly focused on lane-lane topology, i.e., detecting the centerlines to construct a lane graph. Extraction of the online lane topology structure is first proposed by STSU [173], which extended the task to include centreline extraction and lane connectivity relationships. TopoRoad [174] uses a set of directed lane curves and their interactions to represent road topology. Can et al. [175] provide additional supervision of the relationship by considering the centerlines as cluster centers to assign objects. LaneGAP [156] utilizes the shortest path algorithm in graph theory to transform lane topology into a series of overlapping paths and directly obtains information about these complete paths through end-to-end learning. CenterLineDet [176] regard centerlines as vertices and design a graph model to update centerline topology. Recently, lane-traffic topology is additionally introduced by the OpenLaneV2 [153] dataset, further improving the understanding of scene structure. Aiming at a complete and diverse driving scene graph, TopoNet  [177] explicitly models the connectivity of centerlines within the network and incorporates traffic elements into the task. TopoMLP  [154] leverage position embedding [138] to enhance topology modeling. LaneSegNet [155] proposes a unified representation for integrating lanes and centerlines. It introduces a lane attention mechanism to facilitate the learning of topological relationships between centerlines and lanes. To improve lane topology inference, TopoLogic [157] introduces an efficient post-processing that integrates the geometric distance between centerline endpoints and the semantic similarity of lane queries within a high-dimensional space.
@@ VII Future Direction
$\bullet$ General and Unified Lane Modeling. Effectively modeling arbitrarily shaped lanes without compromising efficiency remains a significant challenge. When the scenario extends from a monocular camera’s front view to multi-camera surround views, the presence of numerous U-shaped or nearly horizontal lanes becomes common. In such cases, modeling approaches that heavily rely on prior knowledge, such as grids-based modeling for row-wise classification, or line anchor-based modeling which learns the longitudinal equidistant offset points, are unsuitable. In contrast, mask-based modeling methods are more reliable despite their lower performance and efficiency. The existing vectorized map element modeling methods [170, 148] provide valuable guidance on this issue. Recently, Lane2Seq [141] unified 2D lane detection through sequence generation. This approach also serves as a promising direction for subsequent studies, although its efficiency still requires further improvement.
@@
$\bullet$ Multi-modal Lane Detection. In recent years, LiDAR-based lane detection benchmarks and methods [178, 179, 180, 181, 182, 183, 184] gain attention as another minority approach. Although the 3D information can be directly offered by LiDAR, its shorter perception range and high cost have made camera-based methods more prevalent. However, LiDAR provides the advantages by remaining unaffected by lighting changes and delivering accurate depth information, effectively compensating for the limitations of cameras. The integration of LiDAR and camera data demonstrates significant effectiveness in enhancing performance, which is widely validated in the domain of 3D perception [185, 186, 187, 188]. However, there is currently a lack of dedicated research focusing on multi-modal lane detection [189, 190, 191].
@@
$\bullet$ Label and Model Efficient Lane Detection. The existing lane detection methods mainly focus on supervised learning, which needs a lot of annotations for training, which leads to huge manual costs. Thus, developing annotation-efficient lane detection algorithms is necessary. WS-3D-Lane [192] uses 2D lane labels to weakly supervise 3D lane detection, which is valuable for research and mass production. Furthermore, unsupervised lane detection [193, 194, 195, 196] is also a promising direction, although the related works are less.
@@
$\bullet$ Lane Detection in End-to-End Autonomous Driving. The CVPR Best Paper, UniAD [197], generates significant interest [198, 199, 200, 201, 202] in both academia and industry regarding the development of end-to-end autonomous driving systems. Unlike the conventional modular architecture of ”perception, prediction, and planning,” end-to-end autonomous driving directly outputs vehicle motion planning results from sensor data in a fully differentiable manner. Within this framework, lane detection no longer outputs explicit lane coordinate values but instead functions as a module providing intermediate representations of lanes. However, this approach often encounters challenges such as limited interpretability and inadequate generalization, particularly in complex road scenarios, including curved roads and multi-lane switching. Future research can explore hybrid architectures that incorporate specific lane detection outputs—such as lane centerlines, lane width, and curvature—as prior knowledge into intermediate representations within end-to-end models. This integration can enable the network to better capture the structural characteristics of roads and ensure the preservation of critical lane-level information in the decision-making process.
@@
$\bullet$ Visual Reasoning for Lane Detection. The advent of large language models (LLMs) and vision-language Models (VLMs) unlock significant potential for multimodal artificial intelligence systems to perceive the real world, make decisions, and control tools with a capability akin to human cognition. Recently, LLMs and VLMs are integrated into autonomous driving systems, primarily focusing on visual reasoning related to dynamic obstacles, the generation of future trajectories, and the detailed control signals of ego-vehicles [203, 204, 205]. In contrast, relatively few studies explore visual inference concerning static obstacles, such as lanes. Fortunately, a new benchmark [206] specifically designed for large-scale visual reasoning in understanding maps and traffic scenarios is emerged. By training on extensive traffic scene data, the models can derive insights from complex multimodal driving resources, including map data, traffic regulations, and incident reports. This enables them to enhance vehicle navigation and planning with safety and efficiency parameters, while also adapting to dynamic road conditions with an understanding that closely resembles human intuition.
@@
$\bullet$ Roadside Lane Detection. The current perception capabilities in autonomous driving primarily focus on ego vehicles. While vehicle-based perception systems capture the immediate surrounding environment, their range is limited to short distances. In contrast, roadside cameras, mounted on utility poles several meters above ground, enable remote perception with minimal visual obstructions. Recently, roadside 3D object detection datasets [207, 208] and corresponding methods [209] develop to advance 3D perception tasks in roadside scenes, facilitating potential collaboration between vehicles and infrastructure. However, there are still no established benchmarks or methods specifically for roadside lane detection. Roadside lane detection can effectively substitute for manual monitoring of lane violations or illegal lane changes, offering significant potential for applications in security.
@@ VIII Conclusion
This survey comprehensively reviews the latest progress in monocular lane detection based on deep learning, covering both 2D and 3D lane detection methods in recent years. Four core designs in lane detection algorithms through theoretical analysis and experimental evaluation are identified: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the detection of obscure lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. In addition, this article also reviews extended work on monocular lane detection to help readers have a more comprehensive understanding of the development process of lane detection. Finally, the future research directions for lane detection are pointed out.
@@
$\bullet$ Results on other 2D lane detection datasets. Table VII shows the performance comparison of Tusimple [35], LLAMAS [39] and CurveLanes [40].
TABLE VII: enchmark results of 2D lane detection methods on Tusimple [35], LLAMAS [39] and CurveLanes [40].
@@
TABLE VIII: Benchmark results of 3D lane detection methods on Apollo 3DLane [23].
@@
TABLE IX: Benchmark results of 3D lane detection methods on ONCE-3DLanes [44].
@@
The image coordinate system is employed to represent the perspective projection from the camera coordinate system onto the image plane. When the camera distortion is disregarded, the relationship between a 3D point and its image plane projection can be simplified using a pinhole model. The image coordinates $(x_{i},y_{i})$ are determined by Eqn. 15:
@@
The translation and scaling transformation links the image coordinate framework with the pixel coordinate framework. Let $\alpha$ and $\beta$ denote the scaling factors for the x-axis and y-axis, respectively, while $c_{x}$ and $c_{y}$ represent the translation values shifting the origin of the coordinate system. The pixel coordinates $(u,v)$ can be mathematically formulated as shown in Eqn. 16:
@@
With Eqn. 15 and Eqn. 16, setting $f_{x}=\alpha f$, $f_{y}=\beta f$, we could derive Eqn. 17:

### v3

@@ VI-D Lane Topology Reasoning
Topology reasoning aims to comprehensively understand road scenes and present drivable routes in autonomous driving. It requires detecting road centerlines and traffic elements, further reasoning their topology relationship, i.e., lane-lane topology, and lane-traffic topology. Directly using vehicle-mounted sensors to detect lane topology has become popular due to their practical value.
The early topology reasoning works mainly focus on lane-lane topology, i.e., detecting the centerlines to construct a lane graph. Extraction of the lane topology task is first proposed by STSU [175], which predicts the centreline and lane connectivity relationships. TopoRoad [176] uses a set of directed lane curves and their interactions to represent road topology. Can et al. [177] provide additional supervision of the relationship by considering the centerlines as cluster centers to assign objects. LaneGAP [158] utilizes the shortest path algorithm in graph theory to transform lane topology into a series of overlapping paths and directly obtains information about these complete paths through end-to-end learning. CenterLineDet [178] regards centerlines as vertices and designs a graph model to update centerline topology. Recently, lane-traffic topology is additionally introduced by the OpenLane-V2 [155] dataset, further improving the understanding of scene structure. Aiming at a complete and diverse driving scene graph, TopoNet [179] explicitly models the connectivity of centerlines within the network and incorporates traffic elements into the task. TopoMLP [156] leverages position embedding [139] to enhance topology modeling. LaneSegNet [157] proposes a unified representation for integrating lanes and centerlines. It introduces a lane attention mechanism to facilitate the learning of topological relationships between centerlines and lanes. To improve lane topology inference, TopoLogic [159] introduces an efficient post-processing that integrates the geometric distance between centerline endpoints and the semantic similarity of lane queries within a high-dimensional space.
@@ VII Future Direction
This section outlines potential directions for future research in lane detection. The scope of our discussion includes the improvable issues within the field, underexplored subfields, and relevant tasks outside this area that hold significant research value.
@@
$\bullet$ General and Unified Lane Modeling. Effectively modeling lanes of arbitrary shapes without compromising efficiency remains a significant challenge. When the scenario extends from a monocular camera’s front view to multi-camera surround views, the presence of numerous U-shaped or nearly horizontal lanes becomes common. In such cases, modeling approaches that heavily rely on prior knowledge, such as grids-based modeling for row-wise classification, or line anchor-based modeling for learning the longitudinal equidistant offset points, are unsuitable. In contrast, mask-based modeling methods are more reliable despite their lower performance and efficiency. The existing vectorized map element modeling methods [172, 150] provide valuable guidance on this issue. Recently, Lane2Seq [143] unifies 2D lane detection through sequence generation. This approach also serves as a promising direction for subsequent studies, although its efficiency still requires further improvement.
@@
$\bullet$ Multi-modal Lane Detection. In recent years, LiDAR-based lane detection benchmarks and methods [180, 181, 182, 183, 184, 185, 186] gain attention as another minority approach. Although the 3D information can be directly offered by LiDAR, its shorter perception range and high cost have made camera-based methods more prevalent. However, LiDAR provides the advantages by remaining unaffected by lighting changes and delivering accurate depth information, effectively compensating for the limitations of cameras. The integration of LiDAR and camera data demonstrates significant effectiveness in enhancing performance, which is widely validated in the domain of 3D perception [187, 188, 189, 190]. However, there are notably few dedicated works focusing on multi-modal lane detection [191, 192, 193].
@@
$\bullet$ Label Efficient Lane Detection. The existing lane detection methods mainly focus on supervised learning, which needs a lot of annotations for training, which leads to huge manual costs. Thus, developing annotation-efficient lane detection algorithms is necessary. WS-3D-Lane [194] uses 2D lane labels to weakly supervise 3D lane detection, which is valuable for research and mass production. Furthermore, unsupervised lane detection [195, 196, 197, 198] is also a promising direction, although the related works are limited.
@@
$\bullet$ Lane Detection in End-to-End Autonomous Driving. The CVPR Best Paper, UniAD [199], attracts significant interest [200, 201, 202, 203, 204] in both academia and industry regarding the development of end-to-end autonomous driving systems. Unlike the conventional modular architecture of ’perception-prediction-planning’, end-to-end autonomous driving directly outputs vehicle motion planning results from sensor data in a fully differentiable manner. Within this framework, lane detection no longer outputs explicit lane coordinate values but instead functions as a module providing intermediate representations of lanes. However, this approach often encounters challenges such as limited interpretability and inadequate generalization, particularly in complex road scenarios, including curved roads and multi-lane switching. Future research might explore hybrid architectures that incorporate specific lane detection outputs, such as lane centerlines, lane width, and curvature, as prior knowledge into intermediate representations within end-to-end models. This integration can enable the network to better capture the structural characteristics of roads and ensure the preservation of critical lane-level information in the decision-making process.
@@
$\bullet$ Visual Reasoning for Lane Detection. The advent of large language models (LLMs) and vision-language Models (VLMs) unlock significant potential for multimodal artificial intelligence systems to perceive the real world, make decisions, and control tools with a capability akin to human cognition. Recently, LLMs and VLMs are integrated into autonomous driving systems, primarily focusing on visual reasoning related to dynamic objects, the generation of future trajectories, and the detailed control signals of ego-vehicles [205, 206, 207]. In contrast, relatively few studies explore visual inference concerning static objects, such as lanes. Fortunately, a new benchmark [208] specifically designed for large-scale visual reasoning in understanding maps and traffic scenarios is emerged. By training on extensive traffic scene data, the models can derive insights from complex multi-modal driving resources, including map data, traffic regulations, and incident reports. This enables them to enhance vehicle navigation and planning with safety and efficiency parameters, while also adapting to dynamic road conditions with an understanding that closely resembles human intuition.
@@
$\bullet$ Roadside Lane Detection. The current perception capabilities in autonomous driving primarily focus on ego vehicles. While vehicle-based perception systems capture the immediate surrounding environment, their range is limited to short distances. In contrast, roadside cameras, mounted on utility poles several meters above ground, enable remote perception with minimal visual obstructions. Recently, roadside 3D object detection datasets [209, 210] and corresponding methods [211, 212, 213] are developed to promote 3D perception tasks in roadside scenes, facilitating potential collaboration between vehicles and infrastructure. However, there are still no established benchmarks or methods specifically for roadside lane detection. Roadside lane detection can effectively substitute for manual monitoring of lane violations or illegal lane changes, offering significant potential for applications in security.
@@ VIII Conclusion
This survey comprehensively reviews the latest progress in monocular lane detection based on deep learning, covering both 2D and 3D lane detection methods in recent years. Four core designs in lane detection algorithms are identified through theoretical analysis and experimental evaluation: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the inference on the obscure lanes; (4) Perspective effect elimination, providing accurate 3D lanes for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. In addition, this article also reviews extended works on monocular lane detection to provide readers with a more comprehensive understanding of the development of lane detection. Finally, the future research directions for lane detection are pointed out.
@@
$\bullet$ Results on other 2D lane detection datasets. Table VII shows the performance comparison on Tusimple [35], LLAMAS [39] and CurveLanes [40].
TABLE VII: Benchmark results of representative 2D lane detection methods on Tusimple [35], LLAMAS [39] and CurveLanes [40].
@@
TABLE VIII: Benchmark results of representative 3D lane detection methods on Apollo 3DLane [23].
@@
TABLE IX: Benchmark results of representative 3D lane detection methods on ONCE-3DLanes [44].
@@
The image coordinate system is employed to represent the perspective projection from the camera coordinate system onto the image plane. When the camera distortion is disregarded, the relationship between a 3D point and its image plane projection can be simplified using a pinhole model. The image coordinates $(x_{i},y_{i})$ are determined by Eqn. 14:
@@
The translation and scaling transformation links the image coordinate framework with the pixel coordinate framework. Let $\alpha$ and $\beta$ denote the scaling factors for the x-axis and y-axis, respectively, while $c_{x}$ and $c_{y}$ represent the translation values shifting the origin of the coordinate system. The pixel coordinates $(u,v)$ can be mathematically formulated as shown in Eqn. 15:
@@
With Eqn. 14 and Eqn. 15, setting $f_{x}=\alpha f$, $f_{y}=\beta f$, we could derive Eqn. 16: